Roberto invested some money at 7%, and then invested $2000 more than twice this amount at 9%. His total annual income from the two investments was $2930. How much was invested at 9%?

Answer by checkley77(12844)

.09(2x+2,000)+.07x=2,930
.18x+180+.07x=2,930
.25x=2,930-180
.25x=2,750
x=2,750/.25
x=11,000 ans. for the amount invested @ 7%.
2*11,00+2,000
22,000+2,000=24,000 ans. for the amount invested @ 9%.
Proof:
.09*24,000+.07*11,000=2,930
2,160+770=2,930
2,930=2,930